OpenStudy (anonymous):

Solve for x if: 2 – x = -10 A) -1 B) 1 C) -9 D) 12

OpenStudy (anonymous):

12

OpenStudy (whpalmer4):

2-12 = -10 is correct 2-x=-10 add x to each side 2-x+x = -10+x 2 = -10+x add 10 to each side 2+10 = 10-10+x 12 = x